Advances in Constructive Approximation Nashville, May 14-17, 2003

Support Request Form

Name:

Title/Institute:

Current Position:

Current Adviser (if applicable):

Address:

Email:

Dates Attending Conference:

Arrive: Depart:

Travel:

Total Miles (if you come by car):

Airline: Total Ticket Cost: US$

Hotel:

Name of Hotel:

Sharing Hotel Room with:

Expected Cost of Hotel (your portion if sharing a room): US$

Other support: Who is (partially or fully) paying your expenses?

List all sources and amounts expected:

Total costs you would have to pay from your own pocket without our support^{∗}:

By signing this form, you certify that the above information is complete and accurate: Signature: Date: Social Security Number (if you have one): Type/Class and Number of your Visa for entering the US:

∗ To be able to issue a reimbursement check, you will need to provide us with copies of (1) Passport (frontpage), (2) US Visa, (3) I-94 form (issued at the US border), (4) IAP-66 (if you are on J1 Visa), (5)

Flight ticket. You will also need to fill out a form that you can pick up at the registration desk.